# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ 3.2.2. Convergence of scaled coefficients
$|F(I,\lambda,\mu,\nu,N)|=\left|\frac{\det[\mathrm{e}^{\frac{2i\pi I_{r}\lambda_{s}^{\prime}}{N}}]\det[\mathrm{e}^{\frac{2i\pi I_{r}\mu_{s}^{\prime}}{N}}]\det[\mathrm{e}^{-\frac{2i\pi I_{r}(\nu^{\lor})_{s}^{\prime}}{N}}]}{\Delta(\xi^{I})}\right|\\ \leq\frac{(n!)^{3n}}{|\Delta(\xi^{I})|}.$
On the other hand,
@@
Recall that on $[0,\pi-c]$, one has $\sin(x)\geq\frac{\sin(\pi-c)}{\pi-c}x$. Using Lemma 3.14 for $I$ minimal in its orbit,
@@
$\frac{1}{|\Delta(\xi^{I})|}\leq\left(\frac{N}{c_{n}}\right)^{n(n-1)/2}\prod_{1\leq r<s\leq n}\frac{1}{I_{r}-I_{s}}.$
@@
$\frac{N^{-n+1}}{N^{(n-1)(n-2)/2}}|F(I,\lambda,\mu,\nu,N)|\leq\frac{(n!)^{3n}}{|\Delta(\xi^{I})|}=C_{I}.$
@@
Let $(\alpha,\beta,\gamma)\in\mathcal{H}^{3}$. For $N\geq 1$, let $(\lambda_{N},\mu_{N},\nu_{N})$ be three partition in $\mathbb{Z}_{N-n}^{n}$ such that $|\lambda_{N}|+|\mu_{N}|+|\nu_{N}|=n(N-n)+dN$ for some $d\in\mathbb{Z}_{\geq 0}$ and such that $\frac{1}{N}\lambda_{N}=\alpha+o(1)$, $\frac{1}{N}\mu_{N}=\beta+o(1)$ and $\frac{1}{N}\nu_{N}=\gamma+o(1)$ as $N\rightarrow+\infty$. Let $I=I_{1}>\dots>I_{n-1}>I_{n}=0$ be fixed. Then,
@@
$\lim_{N\rightarrow\infty}N^{-(n-1)(n-2)/2}c_{\lambda_{N},\mu_{N}}^{\nu_{N},d}=\sum_{I:I_{n}=0}\frac{(2\pi)^{-n(n-1)/2}}{\Delta(I)}\det[\mathrm{e}^{2i\pi\alpha_{r}I_{s}}]\det[\mathrm{e}^{2i\pi\beta_{r}I_{s}}]\det[\mathrm{e}^{-2i\pi\gamma_{r}I_{s}}]\\ =J[\gamma|\alpha,\beta].$
where $J[\gamma|\alpha,\beta]$ was defined in (3.27) such that
@@ 4. Puzzles of the quantum cohomology of Grassmannians and their skeleton
4.1. Puzzles and the quantum-LR coefficients
Definition 4.1.
@@ 4.1. Puzzles and the quantum-LR coefficients
The faces of $T_{N}$ are triangles which are called direct (resp. reversed) if the corresponding vertices $(x^{1},x^{2},x^{3})\in V_{n}^{3}$ can be labelled in such a way that $x^{2}-x^{1}=1$ and $x_{3}-x_{1}=e^{i\pi/3}$ (resp. $x_{3}-x_{1}=e^{-i\pi/3}$).
@@
Remark that the height of an edge does not characterize its position, since for example the translations of an edge of type $1$ by $e^{i\pi/3}$ will have the same height.
Let us consider the following set of puzzle pieces, which are considered as the assignment of a label in $\{0,\ldots,7\}$ to edges of $T_{n}$ around a triangular face. Each piece can be rotated by a multiple of $\frac{\pi}{3}$ but not reflected.
@@
The boundary coloring $\partial P$ of a puzzle $P$ is the sequence $(\omega_{0},\omega_{1},\omega_{2})$ such that $\omega_{\ell}$ is the sequence $(P(e))_{e\in\partial_{\ell}T_{N}}$, where $\partial_{\ell}T_{N}$ is the sequence of boundary edges of $T_{N}$ of type $\ell$ ordered by their height.
For any triple $(\omega_{0},\omega_{1},\omega_{2})$ of words in $\{0,1,2\}^{N}$, we denote by $P(\omega_{0},\omega_{1},\omega_{2})$ the set of puzzles whose boundary coloring is $(\omega_{0},\omega_{1},\omega_{2})$.
Based on a conjecture of Knutson and a previous work of Buch [6] relating the quantum cohomology of Grassmannians to the classical cohomology of the two steps flag manifold, it has been proven in [5] that such puzzles describe the quantum Littlewood-Richardson coefficient in the following way.
@@
Let $1\leq n\leq N$ and $\lambda^{0},\lambda^{1},\lambda^{2}$ be partitions of length $n$ with first part smaller than $N-n$ such that $|\lambda^{1}|+|\lambda^{2}|=|\lambda^{0}|+Nd$. Then, $\langle\sigma_{\lambda^{1}},\sigma_{\lambda^{2}},\sigma_{\lambda^{0}}^{\vee}\rangle$ is equal to #$\#P(\omega_{0},\omega_{1},\omega_{2})$ where $\omega_{\ell},\ell\in\{0,1,2\}$ are constructed as follows :
@@
The set of pieces can be further simplified in two steps. First, gluing two pieces with edges having label $2$ along an edge labeled $4,5$ or $6$ yields lozenges with edge labelled $2$ and either $0,1$ or $3$; then concatenating consecutively such lozenges having same label $0,1$ or $3$ and considering also the triangle with all edges labeled $2$ yield the pieces of Figure 10, which are called pieces of type II. Let us then call pieces of type I any piece displayed in Figure 9 which consists of the first three triangles of Figure 8 and pieces obtainged by concatenating two pieces with label $(2,4,1)$ and an arbitrary even number of pieces with label $(4,7,0)$. We will first show that a puzzle $P$ is completely characterize by the position of pieces of type I.
@@ 4.2. Edge, vertex and face partitions
Suppose that $f$ is a triangle of $T_{N}$ with edges $e^{0},e^{1},e^{2}$. Then,
(4.2) $\begin{split}h(e^{0})+h(e^{1})+h(e^{2})=N-1\text{ if $f$ is direct}\\ h(e^{0})+h(e^{1})+h(e^{2})=N-2\text{ if $f$ is reversed}.\end{split}$
@@
Definition 4.6.
@@
The edge partition of $\mathcal{E}$ is the set partitions $\mathcal{P}_{e}$ of $\mathcal{E}$ whose blocks of size greater than one consist of edges of a common type II piece.
@@
  • $e^{0},e^{0^{\prime}}$ are of type $i$ when $f^{1},\ldots,f^{r},e^{0^{\prime}},f^{1^{\prime}},\ldots,f^{r^{\prime}}$ are of type $i+1\mod 3$ for some $i\in\{0,1,2\}$
@@
In the case of a block of order $3$, $\{e^{0},e^{1},e^{2}\}$ for a triangle of $T_{N}$ and the results on the height of the edges is given by Lemma 4.5. The results on the color of the edges is given by the possible coloring of edges of Type I pieces from Figure 9.
@@
Moreover, given a pair $\{e^{1},e^{2}\}$ of edges of same type and same height, one can rephrase the condition of belonging to the strip $S_{B}$ depending on the type of the edge we consider:
  • if $e$ has type $j$, $e\in S_{B}$ if and only if $e_{j}=e^{1}_{j}$ and $e_{j-1}\in[e^{2}_{r-1},e^{1}_{r-1}],$
  • if $e$ has type $j+1$, $e\in S_{B}$ if and only if $e_{j}\in\{e^{1}_{j},e^{1}_{j}+1\}$ and $e_{j-1}\in]e^{2}_{r-1},e^{1}_{r-1}],$
  • if $e$ has type $j-1$, $e\in S_{B}$ if and only if $e_{j}=e^{1}_{j}+1$ and $e_{j-1}\in[e^{2}_{r-1},e^{1}_{r-1}[.$
Lemma 4.9.
@@
Consider two different block $B,B^{\prime}\in\mathcal{P}_{e}$. Let $e\in S_{B}\cap S_{B^{\prime}}$. Remark that any edge of $S_{B}$ or $S_{B^{\prime}}$ is not on the border of the strip if and only if it is neither in $B\cup B^{\prime}$ nor colored $2$. Hence, if $e$ is colored $2$, then $e\in\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$. Suppose by contradiction that $e$ is not colored $2$. Since $B\cap S_{B^{\prime}}=S_{B}\cap B^{\prime}=\emptyset$ and $e$ is not colored $2$, $e$ belongs to the interior of both $S_{B}$ and $S_{B^{\prime}}$ : hence, the two triangular puzzle pieces whose boundary is $e$ belong to $S_{B}$ and $S_{B^{\prime}}$, and we deduce that there is an edge $f^{1}$ colored $0,1$ or $3$ which belong to $S_{B}\cap S_{B^{\prime}}$. If $f^{1}\not\in B\cup B^{\prime}$, $f^{1}$ belong to the interior of $S_{B}$ and $S_{B^{\prime}}$, and thus there exists an edge $f^{2}\in S_{B}\cap S_{B^{\prime}}$ with $f^{2}_{i+1}=f^{1}_{i+1}+1$. Let us repeat the process until there is an edge $f^{s}\in S_{B}\cap S_{B^{\prime}}$ which belongs to either $B$ or $B^{\prime}$. This means that $B\cap S_{B^{\prime}}$ or $B^{\prime}\cap S_{B}$ is not empty, which contradicts the fact proven previously that $S_{B}\cap\mathcal{E}=B$ and $S_{B^{\prime}}\cap\mathcal{E}=B^{\prime}$. Hence, any edge not colored $2$ do not belong to $S_{B}\cap S_{B^{\prime}}$. ∎
@@
If $B=\{e^{1},e^{2}\}$ and $B^{\prime}=\{e^{3},e^{4}\}$ are pairs of the same type $j\in\{0,1,2\}$, the non-crossing condition means that $B\cap S_{B^{\prime}}=\emptyset$ and $B^{\prime}\cap S_{B}=\emptyset$. Suppose that $B$ and $B^{\prime}$ cross, and without loss of generality, assume that $B\cap S_{B^{\prime}}\not=0$. Since $B\cap\partial_{2}S_{B}=\emptyset$, we deduce that $B\cap S_{B^{\prime}}\not\subset\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$. Hence, $S_{B}\cap S_{B^{\prime}}\not\subset\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$.
Reciprocally, suppose that $S_{B}\cap S_{B^{\prime}}\not\subset\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$, and assume without loss of generality that $S_{B}\cap S_{B^{\prime}}\cap\left(S_{B}\setminus\partial_{2}S_{B}\right)\not=\emptyset$. Since $\partial_{2}S_{B}$ is the set of edges of $S_{B}$ of type $j+1$, there exists an edge $e=(x,y)$ of type $j$ or $j-1$ in $S_{B}\cap S_{B^{\prime}}$. If $e$ is of type $j$, this means that $e_{j}=e^{1}_{j}=e^{2}_{j}$ and $e_{j+1}\in[e^{1}_{j+1},e^{2}_{j+1}]$. Similarly, $e_{j}=e^{3}_{j}=e^{4}_{j}$ and $e_{j+1}\in[e^{3}_{j+1},e^{4}_{j+1}]$. Hence, $[e^{1}_{j+1},e^{2}_{j+1}]\cap[e^{3}_{j+1},e^{4}_{j+1}]\not=\emptyset$, and the extremity of one of these interval is contained in the other. Assume without loss of generality that $e^{1}_{j+1}\subset[e^{3}_{j+1},e^{4}_{j+1}]$. Then, $e_{1}$ is an edge of type $j$ such that $e^{1}_{j}=e^{3}_{j}=e^{4}_{j}$ and $e^{1}_{j+1}\in[e^{3}_{j+1},e^{4}_{j+1}]$, thus $e^{1}\in S_{B^{\prime}}$ and thus $B\cap S_{B^{\prime}}\not=\emptyset$. If $e=(x,y)$ is of type $j-1$, then $x_{j-1}=y_{j-1}-1$ and $x_{j}=y_{j}+1$, see (4.1). Hence, the conditions $x_{j},y_{j}\in\{e^{1}_{j},e^{1}_{j}+1\}$ and $x_{j-1},y_{j-1}\in[e^{2}_{j-1},e^{1}_{j-1}]$ from Definition 4.8 yield that $e^{\prime}=(x^{\prime},x)$ with $x^{\prime}_{j}=y_{j}$ and $x^{\prime}_{j-1}=x_{j}$ is an edge of type $j$ which belongs to $S_{B}$. Similarly, $e^{\prime}\in S_{B^{\prime}}$, and the previous reasoning allows to conclude that $B\cap S_{B^{\prime}}\not=\emptyset$ or $B^{\prime}\cap S_{B}\not=\emptyset$.
@@
Suppose that $S_{B}\cap S_{B^{\prime}}\not=\emptyset$. If $B\cap S_{B^{\prime}}\not=\emptyset$ or $B^{\prime}\cap S_{B}\not=\emptyset$, then $B$ and $B^{\prime}$ cross. When $B\cap S_{B^{\prime}}=B^{\prime}\cap S_{B^{\prime}}=\emptyset$, $S_{B}\cap S_{B^{\prime}}\not=\emptyset$ if and only if $\partial_{2}S_{B}\cap S_{B^{\prime}}\not=\emptyset$. One implication is straightforward. For the other implication, remark that there is necessarily an edge $e\in S_{B}\cap S_{B^{\prime}}$ which belongs to $B$ or $\partial_{2}S_{B}$, and since $B\cap S_{B^{\prime}}=\emptyset$, $e\in\partial_{2}S_{B}$. In particular, since $B$ has type $j$, edges of $\partial_{2}S_{B}$ have type $j+1$ and thus $e$ has type $j+1$.
@@
If $x_{j}=e^{1}_{j}$, the latter conditions implies
@@
From the latter lemma, we deduce a description of puzzles in terms of their type I pieces. L
Proposition 4.11.
@@
  • blocks of $\mathcal{P}_{e}$ are either singleton or pairs satisfying the properties of Lemma 4.9,
@@
Let us build the candidate reversed bijection, and consider a subset $\mathcal{E}\subset E_{N}$ with a coloring $c:\mathcal{E}\rightarrow\{0,1,3\}$ and a covering $\mathcal{P}_{v}$ and a partition $\mathcal{P}_{e}$ satisfying the conditions of Proposition 4.11. For each pair $B=(e,e^{\prime})\in\mathcal{P}_{e}$ of type $i$ and color $c$, color all edges of type $i$ (resp. $i+1$, resp. $i+2$) of the strip $S_{B}\setminus\{e,e^{\prime}\}$ with the color $c$ (resp. $2$, resp. $c+5$). Remark that such a coloring is possible, since by properties of Lemma 4.9, any edge belonging to $S_{B}\cap S_{B^{\prime}}$ for two strips of respective types $i,i^{\prime}$ must be included in the boundary $\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$, which consists then of edges of same type $i+1=i^{\prime}+1$ colored $2$ by the above rule.

### v2

@@ 3.2.2. Convergence of scaled coefficients
$|F(I,\lambda,\mu,\nu,N)|=\left|\frac{\det[\mathrm{e}^{\frac{2i\pi I_{r}\lambda_{s}^{\prime}}{N}}]\det[\mathrm{e}^{\frac{2i\pi I_{r}\mu_{s}^{\prime}}{N}}]\det[\mathrm{e}^{-\frac{2i\pi I_{r}(\nu)_{s}^{\prime}}{N}}]}{\Delta(\xi^{I})}\right|\\ \leq\frac{n^{3n}}{|\Delta(\xi^{I})|}.$
First,
@@
Recall that on $[0,c]$ for $0<c<\pi$, one has by concavity $\sin(x)\geq\frac{\sin(c)}{c}x$. Using Lemma 3.14 for $I$ minimal in its orbit,
@@
$\frac{1}{|\Delta(\xi^{I})|}\leq\left(\frac{N}{2c_{n}}\right)^{n(n-1)/2}\prod_{1\leq r<s\leq n}\frac{1}{I_{r}-I_{s}}.$
@@
$\frac{N^{-n+1}}{N^{(n-1)(n-2)/2}}|F(I,\lambda,\mu,\nu,N)|\leq\frac{n^{3n}}{|\Delta(\xi^{I})|}=C_{I}.$
@@
Let $(\alpha,\beta,\gamma)\in\mathcal{H}^{3}$ such that $\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}=\sum_{i=1}^{n}\gamma_{i}+d$ for $d\in\mathbb{N}$. For $N\geq 1$, let $(\lambda_{N},\mu_{N},\nu_{N})$ be three partition in $\mathbb{Z}_{N-n}^{n}$ such that $|\lambda_{N}|+|\mu_{N}|=|\nu_{N}|+dN$ for some $d\in\mathbb{Z}_{\geq 0}$ and such that $\frac{1}{N}\lambda_{N}=\alpha+o(1)$, $\frac{1}{N}\mu_{N}=\beta+o(1)$ and $\frac{1}{N}\nu_{N}=\gamma+o(1)$ as $N\rightarrow+\infty$. Let $I=I_{1}>\dots>I_{n-1}>I_{n}=0$ be fixed. Then,
@@
$\lim_{N\rightarrow\infty}N^{-(n-1)(n-2)/2}c_{\lambda_{N},\mu_{N}}^{\nu_{N},d}=\sum_{I:I_{n}=0}\frac{(2\pi)^{-n(n-1)/2}}{\Delta(I)}\det[\mathrm{e}^{2i\pi\alpha_{r}I_{s}}]\det[\mathrm{e}^{2i\pi\beta_{r}I_{s}}]\det[\mathrm{e}^{-2i\pi\gamma_{r}I_{s}}]\\ =J[\gamma|\alpha,\beta],$
where $J[\gamma|\alpha,\beta]$ was defined in (3.25) and is such that
@@ 4. Puzzles of the quantum cohomology of Grassmannians and their skeleton
4.1. Triangular grid
Definition 4.1 (Triangular grid).
@@ 4.1. Triangular grid
The set $F_{N}$ of faces of $T_{N}$ are triangles which are called direct (resp. reversed) if the corresponding vertices $(x^{1},x^{2},x^{3})\in V_{N}^{3}$ can be labelled in such a way that $x^{2}-x^{1}=1$ and $x_{3}-x_{1}=e^{i\pi/3}$ (resp. $x_{3}-x_{1}=e^{-i\pi/3}$).
Any union of two triangles sharing an edge $e$ is called a lozenge, and $e$ is then called the middle edge of the lozenge.
We denote by $F_{N}^{+}$ (resp. $F_{N}^{-}$) the set of directed (resp. reversed) triangles, so that $F_{N}=F_{N}^{+}\cup F_{N}^{-}$. For $e\in E_{N},f\in F_{N}$, we write $e\in f$ when $e$ is an edge on the boundary of $f$.
@@
We denote by $E_{k}^{(\ell)}$ the set of edges of type $\ell$. Remark that the height of an edge does not characterize its position, since for example the translations of an edge of type $1$ by $e^{i\pi/3}$ will have the same height.
Definition 4.3 (Discrete boundary).
The boundary $\partial T_{N}$ of the triangular grid $T_{N}$ is the set of edges $(x,y)$ lying on the boundary of the triangle $[0,N]\cup[N,Ne^{i\pi/3}]\cup[0,Ne^{i\pi/3}]$.
The boundary $\partial T_{N}$ can be decomposed into three subsets $\partial T_{N}^{(i)}$, $0\leq i\leq 2$, where each set $\partial T_{N}^{(i)}$ consists of edges of type $i$. The coordinates of the corresponding edges are then the following.
$\partial T_{N}^{(0)}$ $=\left((r,N-r,0),0\leq r\leq N-1\right),$
$\partial T_{N}^{(1)}$ $=\left((0,r,N-r),0\leq r\leq N-1\right),$
$\partial T_{N}^{(2)}$ $=\left((N-r,0,r),0\leq r\leq N-1\right).$
4.2. Puzzles and the quantum-LR coefficients
Let us consider the set of puzzle pieces given in Figure 8, which are considered as the assignment of a label in $\{0,\ldots,7\}$ to edges of $T_{N}$ around a triangular face. Each piece can be rotated by a multiple of $\frac{\pi}{3}$ but not reflected.
@@ 4.2. Puzzles and the quantum-LR coefficients
The boundary coloring $\partial P$ of a puzzle $P$ is the sequence $(\omega_{0},\omega_{1},\omega_{2})$ such that $\omega_{\ell}$ is the sequence $(P(e))_{e\in\partial^{(\ell)}T_{N}}$, where $\partial^{(\ell)}T_{N}$ is the sequence of boundary edges of $T_{N}$ of type $\ell$ ordered by their height.
For any triple $(\omega_{0},\omega_{1},\omega_{2})$ of words in $\{0,1,2\}^{N}$, we denote by $P(\omega_{0},\omega_{1},\omega_{2})$ the set of puzzles whose boundary coloring is $(\omega_{0},\omega_{1},\omega_{2})$. For $0\leq k_{0}\leq k_{0}+k_{1}\leq N$, denote by $\mathbb{F}l(k_{0},k_{1},N)$ the two-step flag manifold
$\mathbb{F}l(k_{0},k_{1},N)=\{V_{0}\subset V_{1}\subset\mathbb{C}^{N},\,\dim V_{0}=k_{0},\dim V_{1}=k_{0}+k_{1}\}.$
The cohomology ring $H\mathbb{F}l(k_{0},k_{1},N)$ admits a basis $\{\sigma_{\omega}\}$ of Schubert cycles indexed by words in $\{0,1,2\}^{N}$ with $k_{0}$ occurrences of 00 and $k_{1}$ occurrences of $1$. Proving a conjecture of Knutson, it has been shown in [6] that the previously constructed puzzles describe the structure constants of $H\mathbb{F}l(k_{0},k_{1},N)$.
@@
For any triple $(\omega_{0},\omega_{1},\omega_{2})$ of words in $\{0,1,2\}^{N}$ with same number of occurrences $k_{0}$ of 00 and $k_{1}$ of $1$,
$\langle\sigma_{\omega_{0}}\sigma_{\omega_{1}}\sigma_{\omega_{2}},\sigma_{0}\rangle_{H\mathbb{F}l(k_{0},k_{1},N)}=\#P(\omega_{0},\omega_{1},\omega_{2}),$
where $\sigma_{0}$ is the fundamental class of $H\mathbb{F}l(k_{0},k_{1},N)$.
Thanks to a previous work [7] relating the quantum cohomology of Grassmannians to the classical cohomology of the two-step flag manifold, Theorem 4.5 yields a similar expression in terms of puzzles for the quantum Littlewood-Richardson coefficients.
Corollary 4.6 ([6]).
Let $1\leq n\leq N$ and $\lambda^{0},\lambda^{1},\lambda^{2}$ be partitions of length $n$ with first part smaller than $N-n$ such that $|\lambda^{1}|+|\lambda^{2}|=|\lambda^{0}|+Nd$. Then, $c_{\lambda^{1},\lambda^{2}}^{\lambda^{0},d}=\#P(\omega_{0},\omega_{1},\omega_{2})$, where $\omega_{\ell},\ell\in\{0,1,2\}$ are constructed as follows :
@@
The goal of this section and the next one is then to give a convex formulation of the latter results, yielding Theorem 5.3 for the expression of the structure constants of $H\mathbb{F}l(k_{0},k_{1},N)$ and Corollary 5.4 for the corresponding result concerning the quantum Littlewood-Richardson coefficients.
4.3. Edge, vertex and face partitions
The set of pieces can be further simplified in two steps. First, gluing two pieces with edges having label $2$ along an edge labeled $4,5$ or $6$ yields lozenges with edge labelled $2$ and either $0,1$ or $3$; then concatenating consecutively such lozenges having same label $0,1$ or $3$ and considering also the triangle with all edges labeled $2$ yield the pieces of Figure 10, which are called pieces of type II. Let us then call pieces of type I any piece displayed in Figure 9 which consists of the first three triangles of Figure 8 and pieces obtained by concatenating two pieces with label $(2,4,1)$ and an arbitrary even number of pieces with label $(4,7,0)$. We will first show that a puzzle $P$ is completely characterized by the position of pieces of type I.
@@ 4.3. Edge, vertex and face partitions
Suppose that $f$ is a face of $T_{N}$ with edges $e^{0},e^{1},e^{2}$. Then,
(4.2) $\begin{split}h(e^{0})+h(e^{1})+h(e^{2})&=N-1\text{ if $f$ is direct}\\ h(e^{0})+h(e^{1})+h(e^{2})&=N-2\text{ if $f$ is reversed}.\end{split}$
@@
Definition 4.8 (Edge set, vertex, edge and face partitions).
@@
The edge partition of $\mathcal{E}$ is the set partition $\mathcal{P}_{e}$ of $\mathcal{E}$ whose blocks of size greater than one consist of edges of a common type II piece.
@@
  • $e^{0},e^{0^{\prime}}$ are of type $i$ and $f^{1},\ldots,f^{r},f^{1^{\prime}},\ldots,f^{r^{\prime}}$ are of type $i+1\mod 3$ for some $i\in\{0,1,2\}$,
@@
In the case of a block of order $3$, $\{e^{0},e^{1},e^{2}\}$ is a triangle of $T_{N}$ and the results on the height of the edges is given by Lemma 4.7. The results on the color of the edges is given by the possible coloring of edges of Type I pieces from Figure 9.
@@
Moreover, given a pair $B=\{e^{1},e^{2}\}$ of edges of same type $j$ and same height, one can rephrase the condition of belonging to the strip $S_{B}$ depending on the type of the edge we consider:
  • if $e$ has type $j$, $e\in S_{B}$ if and only if $e_{j}=e^{1}_{j}$ and $e_{j-1}\in[e^{2}_{j-1},e^{1}_{j-1}],$
  • if $e$ has type $j+1$, $e\in S_{B}$ if and only if $e_{j}\in\{e^{1}_{j},e^{1}_{j}+1\}$ and $e_{j-1}\in]e^{2}_{j-1},e^{1}_{j-1}],$
  • if $e$ has type $j-1$, $e\in S_{B}$ if and only if $e_{j}=e^{1}_{j}+1$ and $e_{j-1}\in[e^{2}_{j-1},e^{1}_{j-1}[.$
Lemma 4.11 (Edge partition pairs are admissible).
@@
Consider two different blocks $B,B^{\prime}\in\mathcal{P}_{e}$. Let $e\in S_{B}\cap S_{B^{\prime}}$. Remark that any edge of $S_{B}$ or $S_{B^{\prime}}$ is not on the border of the strip if and only if it is neither in $B\cup B^{\prime}$ nor colored $2$. Hence, if $e$ is colored $2$, then $e\in\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$. Suppose by contradiction that $e$ is not colored $2$. Since $B\cap S_{B^{\prime}}=S_{B}\cap B^{\prime}=\emptyset$ and $e$ is not colored $2$, $e$ belongs to the interior of both $S_{B}$ and $S_{B^{\prime}}$ : hence, the two triangular puzzle pieces whose boundary is $e$ belong to $S_{B}$ and $S_{B^{\prime}}$, and we deduce that there is an edge $f^{1}$ colored $0,1$ or $3$ which belong to $S_{B}\cap S_{B^{\prime}}$. If $f^{1}\not\in B\cup B^{\prime}$, $f^{1}$ belong to the interior of $S_{B}$ and $S_{B^{\prime}}$, and thus there exists an edge $f^{2}\in S_{B}\cap S_{B^{\prime}}$ with $f^{2}_{i+1}=f^{1}_{i+1}+1$. Let us repeat the process until there is an edge $f^{s}\in S_{B}\cap S_{B^{\prime}}$ which belongs to either $B$ or $B^{\prime}$. This means that $B\cap S_{B^{\prime}}$ or $B^{\prime}\cap S_{B}$ is not empty, which contradicts the fact proven previously that $S_{B}\cap\mathcal{E}=B$ and $S_{B^{\prime}}\cap\mathcal{E}=B^{\prime}$. Hence, any edge not colored $2$ does not belong to $S_{B}\cap S_{B^{\prime}}$. ∎
Definition 4.12 (Crossing pairs).
@@
If $B=\{e^{1},e^{2}\}$ and $B^{\prime}=\{e^{3},e^{4}\}$ are pairs of the same type $j\in\{0,1,2\}$, the non-crossing condition means that $B\cap S_{B^{\prime}}=\emptyset$ and $B^{\prime}\cap S_{B}=\emptyset$. Suppose that $B$ and $B^{\prime}$ cross, and without loss of generality, assume that $B\cap S_{B^{\prime}}\not=\emptyset$. Since $B\cap\partial_{2}S_{B}=\emptyset$, we deduce that $B\cap S_{B^{\prime}}\not\subset\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$. Hence, $S_{B}\cap S_{B^{\prime}}\not\subset\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$.
Reciprocally, suppose that $S_{B}\cap S_{B^{\prime}}\not\subset\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$, and assume without loss of generality that $S_{B}\cap S_{B^{\prime}}\cap\left(S_{B}\setminus\partial_{2}S_{B}\right)\not=\emptyset$. Since $\partial_{2}S_{B}$ is the set of edges of $S_{B}$ of type $j+1$, there exists an edge $e=(x,y)$ of type $j$ or $j-1$ in $S_{B}\cap S_{B^{\prime}}$. If $e$ is of type $j$, this means that $e_{j}=e^{1}_{j}=e^{2}_{j}$ and $e_{j+1}\in[e^{1}_{j+1},e^{2}_{j+1}]$. Similarly, $e_{j}=e^{3}_{j}=e^{4}_{j}$ and $e_{j+1}\in[e^{3}_{j+1},e^{4}_{j+1}]$. Hence, $[e^{1}_{j+1},e^{2}_{j+1}]\cap[e^{3}_{j+1},e^{4}_{j+1}]\not=\emptyset$, and the extremity of one of these intervals is contained in the other. Assume without loss of generality that $e^{1}_{j+1}\subset[e^{3}_{j+1},e^{4}_{j+1}]$. Then, $e_{1}$ is an edge of type $j$ such that $e^{1}_{j}=e^{3}_{j}=e^{4}_{j}$ and $e^{1}_{j+1}\in[e^{3}_{j+1},e^{4}_{j+1}]$, thus $e^{1}\in S_{B^{\prime}}$ and thus $B\cap S_{B^{\prime}}\not=\emptyset$. If $e=(x,y)$ is of type $j-1$, then $x_{j-1}=y_{j-1}-1$ and $x_{j}=y_{j}+1$, see (4.1). Hence, the conditions $x_{j},y_{j}\in\{e^{1}_{j},e^{1}_{j}+1\}$ and $x_{j-1},y_{j-1}\in[e^{2}_{j-1},e^{1}_{j-1}]$ from Definition 4.10 yield that $e^{\prime}=(x^{\prime},x)$ with $x^{\prime}_{j}=y_{j}$ and $x^{\prime}_{j-1}=x_{j}$ is an edge of type $j$ which belongs to $S_{B}$. Similarly, $e^{\prime}\in S_{B^{\prime}}$, and the previous reasoning allows to conclude that $B\cap S_{B^{\prime}}\not=\emptyset$ or $B^{\prime}\cap S_{B}\not=\emptyset$.
@@
Suppose that $S_{B}\cap S_{B^{\prime}}\not=\emptyset$. If $B\cap S_{B^{\prime}}\not=\emptyset$ or $B^{\prime}\cap S_{B}\not=\emptyset$, then $B$ and $B^{\prime}$ cross. When $B\cap S_{B^{\prime}}=B^{\prime}\cap S_{B}=\emptyset$, then $S_{B}\cap S_{B^{\prime}}\not=\emptyset$ if and only if $\partial_{2}S_{B}\cap S_{B^{\prime}}\not=\emptyset$. One implication is straightforward. For the other implication, remark that there is necessarily an edge $e\in S_{B}\cap S_{B^{\prime}}$ which belongs to $B$ or $\partial_{2}S_{B}$, and since $B\cap S_{B^{\prime}}=\emptyset$, $e\in\partial_{2}S_{B}$. In particular, since $B$ has type $j$, edges of $\partial_{2}S_{B}$ have type $j+1$ and thus $e$ has type $j+1$.
@@
If $x_{j}=e^{1}_{j}$, the latter conditions imply
@@
From the latter lemma, we deduce a description of puzzles in terms of their type I pieces.
Proposition 4.14 (Partitions determine puzzles).
@@
  • blocks of $\mathcal{P}_{e}$ are either singletons or pairs satisfying the properties of Lemma 4.11,
@@
Let us build the candidate inverse bijection, and consider a subset $\mathcal{E}\subset E_{N}$ with a coloring $c:\mathcal{E}\rightarrow\{0,1,3\}$ and a covering $\mathcal{P}_{v}$ and a partition $\mathcal{P}_{e}$ satisfying the conditions of Proposition 4.14. For each pair $B=(e,e^{\prime})\in\mathcal{P}_{e}$ of type $i$ and color $c$, color all edges of type $i$ (resp. $i+1$, resp. $i+2$) of the strip $S_{B}\setminus\{e,e^{\prime}\}$ with the color $c$ (resp. $2$, resp. $c+5$). Remark that such a coloring is possible, since by properties of Lemma 4.11, any edge belonging to $S_{B}\cap S_{B^{\prime}}$ for two strips of respective types $i,i^{\prime}$ must be included in the boundary $\partial_{2}S_{B}\cap\partial_{2}S_{B^{\prime}}$, which consists then of edges of same type $i+1=i^{\prime}+1$ colored $2$ by the above rule.